OpenStudy (anonymous):

In a closed system, Glider A with a mass of 0.40 kg and a speed of 2.00 m/s collides with Glider B at rest with a mass of 1.20 kg. The two interlock and move off. What speed are they moving? A. 0.50 m/s B. 1.0 m/s C. 2.0 m/s D. 4.0 m/s

OpenStudy (michele_laino):

here we have to apply the total momentu conservation law. So we can write this: \[\Large {m_1}{v_1} + {m_2}{v_2} = \left( {{m_1} + {m_2}} \right)V\]

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

where V is the requested speed: \[\Large V = \frac{{{m_1}{v_1} + {m_2}{v_2}}}{{{m_1} + {m_2}}} = \frac{{0.4 \times 2 + 1.2 \times 0}}{{0.4 + 1.2}} = ...m/\sec \]

OpenStudy (michele_laino):

\[\large V = \frac{{{m_1}{v_1} + {m_2}{v_2}}}{{{m_1} + {m_2}}} = \frac{{0.4 \times 2 + 1.2 \times 0}}{{0.4 + 1.2}} = ...m/\sec \]

OpenStudy (anonymous):

so we get 0.5? so choice A is our solution?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay!! thanks!!:)

OpenStudy (michele_laino):

:)